# The vanishing of the first tight Hilbert coefficient for Buchsbaum rings

### v1

@@ Abstract.
Key words and phrases:
tight Hilbert coefficients, Buchsbaum rings
2020 Mathematics Subject Classification: 13A35, 13H10, 13D40, 13D45
@@ 1. Introduction
This conjecture is true for Cohen-Macaulay local rings in [4, Theorem 4.4], and tight Buchsbaum local rings in [2, Corollary 4.6]. We will prove this conjecture for Buchsbaum local rings in Theorem 3.4. We also prove that the condition $(S_{2})$ of the ring cannot be omitted in Proposition 3.7.
@@ Acknowledgement
This article was written during many visits of the third author to Vietnam Institute for Advanced Studies in Mathematics. He thanks sincerely the institute for their hospitality and valuable supports. The third author also thanks to Basque Center for Applied Mathematics for their partial support. He is grateful to Kriti Goel, Linquan Ma and Ilya Smirnov for their comments on the paper.
@@ 2.1. Hilbert coefficients
Let $Q$ be a parameter ideal of $R$. Then $e_{1}(Q)\leq 0$. Moreover, $R$ is Cohen-Macaulay if and only if $e_{1}(Q)=0$ for some $Q$ provided $R$ is unmixed.
@@
Let $(R,\mathfrak{m})$ be a local ring such that $\widehat{R}$ is reduced and equidimensional. Then for all parameter ideals $Q$ we have $\overline{e}_{1}(Q)\geq 0$. Moreover, if $\overline{e}_{1}(Q)=0$ for some parameter ideal $Q$, then $R$ is regular and $\nu(\mathfrak{m}/Q)\leq 1$ provided $\widehat{R}$ is reduced and $(S_{2})$.
@@
Let $(R,\mathfrak{m})$ be a local ring of prime characteristic $p$ such that $\widehat{R}$ is reduced and equidimensional. Then for all parameter ideals $Q$ we have $e^{*}_{1}(Q)\geq 0$. Moreover, if $e^{*}_{1}(Q)=0$ for some parameter ideal $Q$, then $R$ is $F$-rational provided $R$ is Cohen-Macaulay.
@@ 2.2. Buchsbaum and tight Buchsbaum rings
We recall definition of Buchsbaum rings.
@@
Let $(R,\mathfrak{m})$ be a Buchsbaum local ring of dimension $d$ and $Q$ any parameter ideal. We have
@@
Hilbert function of parameter ideals in Buchsbaum rings can be understood in terms of local cohomology, see [13, Corollary 3.2], [14, Corollary 4.2].
@@ 3. The main results
Since $Q^{n}Q^{*}\subseteq(Q^{n+1})^{*}$ and by Lemma 3.1 so
@@
Since $R$ is Buchsbaum, by Theorem 2.6 for all parameter ideal $Q$ we have
@@
We next show that the $(S_{2})$ condition can not be removed in Theorem 3.4. First, we recall some facts about $S_{2}$-ification of the ring by [6]. Suppose $R$ is an unmixed local ring. We shall say that a ring $S$ is an $S_{2}$-ification of $R$ if it lies between $R$ and its total quotient ring, is module-finite over $R$, is $(S_{2})$ as an $R$-module, and has the property that for every element $s\in S\setminus R$, the ideal $D(s)$, defined as $\{r\in R\mid rs\in R\}$, has height at least two. If $(R,\mathfrak{m})$ is an unmixed ring then $R$ has $S_{2}$-ification.
@@
We provide some concrete examples.
@@
Hence, the $(S_{2})$ condition is important for Lemma 3.3 and Theorem 3.4.
@@
The main result of this paper should be true for mixed characteristic when we replace $e_{1}^{*}(Q)$ by $e_{1}^{B}(Q)$ for any big Cohen-Macaulay ring $B$ of $R$. We will consider this topic in the future.

### v2

@@ Abstract.
2020 Mathematics Subject Classification: 13A35, 13H10, 13D40, 13D45.
Keywords: tight Hilbert coefficients, Buchsbaum rings.
Corresponding author: Pham Hung Quy
This work is partially supported by a fund of Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 101.04-2023.08.
@@ 1. Introduction
This conjecture is true for Cohen-Macaulay local rings in [4, Theorem 4.4], and tight Buchsbaum local rings in [2, Corollary 4.6]. This paper proves the conjecture under new assumptions; the main theorem is as follows.
Theorem 1.2.
Conjecture 1 holds if we further assume that R is reduced and Buchsbaum.
We also prove that the condition $(S_{2})$ of the ring cannot be omitted in Proposition 3.7.
@@ Acknowledgement
The authors are grateful to the referee for carefully reading the paper. This article was written during many visits of the third author to Vietnam Institute for Advanced Studies in Mathematics. He thanks sincerely the institute for their hospitality and valuable supports. The third author also thanks to Basque Center for Applied Mathematics for their partial support. He is grateful to Kriti Goel, Linquan Ma and Ilya Smirnov for their comments on the paper.
@@ 2.1. Hilbert coefficients
Let $Q$ be a parameter ideal of $R$. Then $e_{1}(Q)\leq 0$. Moreover, assume $R$ is unmixed, $R$ is Cohen-Macaulay if and only if $e_{1}(Q)=0$ for some $Q$.
@@
Let $(R,\mathfrak{m})$ be a local ring such that $\widehat{R}$ is reduced and equidimensional. Then for all parameter ideals $Q$ we have $\overline{e}_{1}(Q)\geq 0$. Moreover, if $\overline{e}_{1}(Q)=0$ for some parameter ideal $Q$, then $R$ is regular and $\nu(\mathfrak{m}/Q)\leq 1$ provided $\widehat{R}$ is $(S_{2})$.
@@
Let $(R,\mathfrak{m})$ be a local ring of prime characteristic $p$ such that $\widehat{R}$ is reduced and equidimensional. Then for all parameter ideals $Q$ we have $e^{*}_{1}(Q)\geq 0$. Assume $R$ is Cohen-Macaulay, then $e^{*}_{1}(Q)=0$ for some parameter ideal $Q$ if and only if $R$ is $F$-rational.
@@ 2.2. Buchsbaum and tight Buchsbaum rings
We recall the definition of Buchsbaum rings.
@@
Let $(R,\mathfrak{m})$ be a Buchsbaum local ring of dimension $d$ and $Q$ any parameter ideal (for example, see [14]). We have
@@
The Hilbert function of parameter ideals in Buchsbaum rings can be understood in terms of local cohomology, see [13, Corollary 3.2], [14, Corollary 4.2].
@@
The distinction between Buchsbaum rings and tight Buchsbaum rings is subtle, and there exist Buchsbaum (or even Cohen-Macaulay) rings that fail to be tight Buchsbaum. For example, if $R=\mathbb{F}_{p}[[X,Y,Z]]/(X^{4}+Y^{4}+Z^{4})$ then it is a Cohen-Macaulay ring. However, $R$ is not tight Buchsbaum since, in this case, the tight closure of the zero submodule of the top local cohomology is not a $\mathbb{F}_{p}$-vector space.
@@ 3. The main results
Since $Q^{n}Q^{*}\subseteq(Q^{n+1})^{*}$ and by Lemma 3.1 we have
@@
Since $R$ is Buchsbaum, by Theorem 2.6 for all parameter ideals $Q$ we have
@@
We next show that the $(S_{2})$ condition can not be removed in Theorem 3.4. First, we recall some facts about $S_{2}$-ification of the ring by [6]. Suppose $R$ is an unmixed local ring. We shall say that a ring $S$ is an $S_{2}$-ification of $R$ if it lies between $R$ and its total quotient ring, is module-finite over $R$, is $(S_{2})$ as an $R$-module, and has the property that for every element $s\in S\setminus R$, the ideal $D(s)$, defined as $\{r\in R\mid rs\in R\}$, has height at least two. If $(R,\mathfrak{m})$ is an unmixed ring then $R$ has a $S_{2}$-ification and this is unique (up to an isomorphism).
Recall that let $R,\mathfrak{m}$ be a Noetherian local ring and $I$ be an ideal of $R$. An element $x\in I$ is called superficial for $I$ if there exists an integer $c$ such that for all $n\geq c$, the following equality holds:
$(I^{n+1}:x)\cap I^{c}=I^{n}.$
@@
We provide some concrete examples. The first example is closely related to Proposition 5.3 of [2].
@@
Hence, the $(S_{2})$ condition is important for Lemma 3.3, and hence for Theorem 3.4.
@@
The main results of this paper may hold in mixed characteristic when $e_{1}^{*}(Q)$ is replaced with $e_{1}^{B}(Q)$ for any big Cohen-Macaulay ring $B$ of $R$. Additionally, it would be interesting to study the analogue of tight Buchsbaum rings in mixed characteristic which we refer to as BCM-Buchsbaum rings.